OpenStudy (anonymous):

A recursive function is shown below: f(1) = 2 and f(n) = f(n - 1) - 5; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 2, 3, 8, 13, 18, ... 2, 7, 12, 17, 22, ... 2, -3, -8, -13, -18, ... 2, -7, -12, -17, -22, ...

OpenStudy (anonymous):

@blahhedup101

OpenStudy (aum):

f(1) = 2 f(n) = f(n - 1) - 5; n > 1 put n = 2 f(2) = f(1) - 5 = 2 - 5 = -3 put n = 3 ....

OpenStudy (aum):

Just the first two terms is sufficient to answer this problem. But it is worth trying a few n values for learning purposes.

OpenStudy (anonymous):

So what's the answer?

OpenStudy (anonymous):

I didn't see this, my bad.

OpenStudy (anonymous):

is it 2,3... or 2, -3.....

OpenStudy (aum):

2, -3, ...

OpenStudy (aum):

It is C or the third choice.